OpenStudy (anonymous):

What is the radius of a circle in which a 30° arc is 2pi inches long? a) 24 inches b) 12 inches c) 2√3 inches

OpenStudy (anonymous):

WILL MEDAL

OpenStudy (danjs):

30/360 is 2pi so the whole circumference is ...

OpenStudy (danjs):

\[12\frac{ 30 }{ 360 } = 360/360 = 12*2\pi = Full ~ Circumference \]

OpenStudy (anonymous):

75.36?

OpenStudy (danjs):

That is the circumference - 24pi , you want the radius C = 2*pi*r 24pi = 2pi*r

OpenStudy (anonymous):

So the radius is 4

OpenStudy (danjs):

\[r = \frac{ 24\pi }{ 2\pi }\]

OpenStudy (anonymous):

69.08

OpenStudy (anonymous):

Oh so the radius is 69.08?

OpenStudy (danjs):

that is not an answer choice... the correct one is there

OpenStudy (anonymous):

C)

OpenStudy (anonymous):

?

OpenStudy (danjs):

\[r = \frac{ 24\pi }{ 2\pi } = \frac{ 24 }{ 2 } *\frac{ \pi }{ \pi } = \frac{ 24 }{ 2 }*1\]

OpenStudy (anonymous):

12?

OpenStudy (anonymous):

So it's b)

OpenStudy (anonymous):

Thanks so much